OpenStudy (anonymous):

HELP:State whether the ratios form a proportion. 16 to 18 and 24 to 27 A. Yes B. No

OpenStudy (jadeishere):

well, simplified, 16/18 = 8/9 and 24/27 = 8/9 so yes, they form a proportion

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Thanks everyone :)

OpenStudy (jadeishere):

yep :)